What is the S?- concentration of a 103M HCI solution which is saturated with H2S ? H,S(aq) + 2H,O Ka = 1.1 x 10-21 2H;0* (aq) + S² (aq). Please give m answered solution. Fast


1
Expert's answer
2021-03-10T06:54:23-0500

HCl "\to" H+ + Cl-

______1.03M ____

H2S "\to" 2H+ + S2-

______1.03___x

[H+]2 [S2-] = Ka

[1.03]2 [x] = 1.1×10-21

x = 1.1×10-21 /(1.03×1.03)

x = 1.03×10-21